OpenStudy (anonymous):

wire of resistance 5 ohm is stretched to twice its original length. What is the reaistance of wire now ? 10 ohm 15 ohm 25 ohm 20 ohm can not be determined

OpenStudy (ujjwal):

\[R=\frac{\rho l}{A}\]Multiplying numerator and denominator by l (length)\[R=\frac{\rho l^2}{V}\]where V is volume Now, let \(R_1\) and \(R_2\) be the resistances so we have \[\frac{R_1}{R_2}=\frac{l^2}{4l^2}\]Here \(R_1\)=5\(\Omega\) So, \(R_2\)=20\(\Omega\)

OpenStudy (anonymous):

Resistance=pl/a ( p is resistivity, l=lenght of wire, a= area of crossection) if L=2l then new resistance =pL/A ( pie*r*r*l=pie*R*R*(2l) new radius of wire, R=r/root2 new area after streching, A=pie*r*r/2 ) new resitance = p(2l)/A substituting value of A, we get new resistance= 4times the old resistance new resistance =4*5ohms = 20 ohms

OpenStudy (jamesj):

Nicely done ujjwal

OpenStudy (ujjwal):

thanks @JamesJ